The perimeter of a rectangle is 72 meters. If the width is 8 meters less than the length, then what is the width of the rectangle?

Found 2 solutions by rfer, checkley77:
Answer by rfer(10432)

2x+2(x+8)=72
2x+2x+16=72
4x=56
x=14=width
x+8=22=length

Answer by checkley77(12569)

W=L-8
2L+2W=72
2L+2(L-8)=72
2L+2L-16=72
4L=72+16
4L=88
L=88/4
L=22 ANS.
W=22-8
W=14 ANS.
PROOF:
2*22+2*14=72
44+28=72
72=72